# Simple Methods Defend RAG Systems Well Against Real-World Attacks

### v1

@@ Abstract
Ensuring safety and in-domain responses for Retrieval-Augmented Generation (RAG) systems is paramount in safety-critical applications, yet remains a significant challenge. To address this, we evaluate four methodologies for Out-Of-Domain (OOD) query detection: GPT-4o, regression-based, Principal Component Analysis (PCA)-based, and Neural Collapse (NC), to ensure the RAG system only responds to queries confined to the system’s knowledge base. Specifically, our evaluation explores two novel dimensionality reduction and feature separation strategies: PCA, where top components are selected using explained variance or OOD separability, and an adaptation of Neural Collapse Feature Separation. We validate our approach on standard datasets (StackExchange and MSMARCO) and real-world applications (Substance Use and COVID-19), including tests against LLM-simulated and actual attacks on a COVID-19 vaccine chatbot. Through human and LLM-based evaluations of response correctness and relevance, we confirm that an external OOD detector is crucial for maintaining response relevance. [^1] https://github.com/toastedqu/rag_safety_pca
Introduction
@@
In high-stakes domains, the accuracy and domain relevance of responses provided by Retrieval-Augmented Generation (RAG) systems are critical for ensuring safety and reliability. One significant challenge these systems face is detecting and handling out-of-domain (OOD) queries, which can compromise performance and safety. For instance, in the medical field, a RAG system for clinical decision support must accurately discern relevant medical information (Giorgi et al. 2024). A failure to do so — such as treating an OOD query about a rare medical condition as if it were in-domain (ID), could result in incorrect medical advice, posing adverse health outcomes. An example is illustrated in Fig. 1 for the COVID-19 domain that shows how an OOD query (right side) can bypass a system’s safeguards and produce a potentially malicious response.  [^2] We treat malicious queries as a subset of out-of-domain queries. This is justified theoretically because our detection methods are trained not merely on KB similarity, but also on broader ID and OOD distributions as described later in the methodology.
While research in RAG predates the era of Large Language Models (LLMs) (Lewis et al. 2020; Guu et al. 2020; Izacard et al. 2023; Ram et al. 2023; Li et al. 2022), early studies generally assumed that input queries were relevant to the available knowledge bases (KB) with no rigorous verification. The rise of LLMs has spurred efforts to enhance RAG systems (Jiang et al. 2023; Cheng et al. 2024; Zakka et al. 2024; Feng et al. 2024), but their integration has heightened safety and trust concerns in both retrieval and generation. The retrieval module’s performance largely depends on embedding quality and semantic search capabilities, which often lag behind lexical search in terms of precision and robustness to keyword variations (Barnett et al. 2024; Reichman and Heck 2024; Chandrasekaran and Mago 2021). Meanwhile, the generation module is susceptible to producing ‘hallucinated’ content that may not be grounded in the source material, leading to misinformation if not properly checked (Bang et al. 2023; Xu, Jain, and Kankanhalli 2024; Zhang et al. 2023). Hence, ensuring domain relevance of queries is crucial for maintaining the reliability and safety of RAG systems.
Nowadays, practitioners commonly attempt to mitigate OOD queries either by relying on the built-in guardrails of LLMs or by issuing additional LLM calls to explicitly judge whether a query is in-domain (Peng et al. 2025). While these methods can improve safety and accuracy, they often introduce significant overhead in terms of latency and API cost. Moreover, LLM guardrails are primarily designed to detect toxic or unsafe content and not domain irrelevant, and thus offer limited protection against benign but unanswerable queries (Dong et al. 2024). In this work, we propose lightweight and interpretable OOD detection techniques that match these baseline approaches while substantially reducing computational cost and inference latency.
[…]
To complement this, we adapt a Neural Collapse (NC)-based feature separation method, originally developed for computer vision, to the textual OOD detection setting. Our novel approach introduces a clustering-based pseudo-labeling step for ID queries, followed by fine-tuning to encourage separation between ID and OOD samples. This method offers a more expressive alternative to linear projection and allows us to evaluate whether higher-capacity representations offer more gains in OOD discrimination, particularly in domains where linear methods may fall short.
In the second part of our study, we extend our evaluation beyond classification by assessing the impact of OOD detection on chatbot responses. Specifically, we examine the relevance and the correctness of chatbot responses.
Our key contributions are as follows:
  • We propose and demonstrate the efficiency of simple methods, including PCA-based and a novel adaptation of NC feature separation, for OOD query detection in four datasets containing 16 different domains
  • Show that simple methods such as PCA and logistic regression are often superior to sophisticated methods like NC and GPT-based classification in these real-world safety-critical scenarios
  • We empirically demonstrate the necessity of an external out-of-domain detector to maintain response relevance, as modern LLM guardrails alone fail to adequately filter out harmful and irrelevant queries
Related Work
Safety Concerns in RAG systems
@@
RAG systems rely on retrieving relevant documents to ensure accurate and trustworthy responses (Lewis et al. 2020), yet incorrect retrievals can severely degrade generation quality (Creswell, Shanahan, and Higgins 2022; Barnett et al. 2024). Efforts to improve retrieval include incorporating topical context (Ahn et al. 2022), conversation history (Shuster et al. 2021), and predictive sentence generation (FLARE) (Jiang et al. 2023). Interestingly, unrelated documents may sometimes enhance generation, while highly ranked but irrelevant ones can harm it (Cuconasu et al. 2024). To mitigate such issues, methods have proposed response skeletons (Cai et al. 2019), prompt-based validation (Yu et al. 2023), Natural Language Inference (NLI) filtering (Yoran et al. 2023), and dynamic reliance on parametric vs. retrieved knowledge (Li et al. 2023; Longpre et al. 2021; Mallen et al. 2023). Our work adds to this literature by introducing new methods for selecting when to answer and verifying the effect of out-of-domain questions.
Adversarial Attacks
@@
Adversarial attacks mislead models through crafted inputs (Zhang et al. 2020), with recent work targeting LLMs to produce harmful content (Zou et al. 2023). For RAG systems, attacks often involve malicious documents that degrade retrieval or generation (Cho et al. 2024; Xue et al. 2024; Shafran, Schuster, and Shmatikov 2024). While early attacks required specific trigger queries (Zou et al. 2024), newer methods exploit query-agnostic poisoning (Chaudhari et al. 2024). Our work mitigates query-based adversarial attacks by detecting when a modified question lacks an answer in the database.
Null-space Projections
Several recent methods leverage null-space projections to enhance OOD detection in ways compatible with Neural Collapse. ViM (Wang et al. 2022) and NuSA (Cook, Zare, and Gader 2007) decompose features into ID-aligned and orthogonal components, using null-space norms as OOD scores. NECO (Ammar et al. 2023) and RankFeat (Song, Sebe, and Wang 2022) exploit the low-rank structure of NC by scoring deviations from class-aligned subspaces. Others use PCA residuals (Ndiour, Ahuja, and Tickoo 2020) or enforce orthogonality between ID and auxiliary OOD features (Wu et al. 2024), further improving separability. These techniques complement our NC-based feature projection and can be integrated into RAG pipelines.
Methods
Out-Of-Domain Detection
Low-Dimensional Methods
[…]
@@ Methods
  • Explained Variance (EVR): In this approach, the final set of PCs consists of the $m$ components with the highest explained variance, where $m\leq k$. This ensures that we retain only the components contributing the most to the dataset’s variance.
[…]
@@
Both criteria for selecting $m$ principal components aim to retain the most informative aspects of the embeddings while reducing dimensionality. This step not only preserves the discriminative power of the embeddings but also enhances computational efficiency for subsequent tasks.
[…]
@@
The semantic-search algorithms operate by mapping the query embeddings from the training set into a $m$-dimensional space derived from the principal component selection process. During inference, these algorithms employ distinct geometric criteria to make a classification decision for a test query $u$ with projected embedding $\mathbf{e}_{u}^{\prime}$. The three algorithms are as follows:
@@
In addition to the semantic-search algorithms, we leverage three simple yet effective machine learning models. These models are trained on the entire training set, which includes both ID and OOD queries, for a binary classification task. During inference, the algorithms classify the test query $u$ with its projected embedding $\mathbf{e}_{u}$ into one of the two classes. We use the following models: a Logistic Regression (LogReg) (Kleinbaum et al. 2002), Support Vector Machines (SVM) (Hearst et al. 1998), and Gaussian Mixture Models (GMM) (Reynolds et al. 2009).
Neural Collapse Feature Separation
For this technique, we adapt the approach of Wu et al. (2024), (originally developed for computer vision) to the domain of text-based OOD detection. To address the lack of explicit class labels in our ID queries, we introduce a clustering-based workaround, followed by fine-tuning and inference steps that mirror the method’s intent.
Clustering.
Since the original method relies on ID samples from clearly labeled classes (e.g., CIFAR10 categories), we must simulate a similar setting for textual queries. To do so, we apply clustering to the ID questions using the KMeans algorithm on BERT-based query embeddings. Each query is then assigned a pseudo-label corresponding to its cluster. These pseudo-labels are used to construct the fully-connected (FC) layer of the classifier, enabling the application of the NC property to text embeddings.
Fine-Tuning.
In this stage, we fine-tune a pretrained Language Model $f_{\theta}$ using a combined batch of ID and auxiliary OOD queries. Let $z_{\text{ID}}\in\mathbb{R}^{d}$ be the last-layer embedding of an ID query and let $w_{y}$ be the normalized weight vector of the FC layer corresponding to its cluster label $y$. We encourage the NC phenomenon by minimizing the following clustering loss, which aligns the ID feature to its respective FC weight:
$\mathcal{L}_{\text{Clu}}=-z_{\text{ID}}^{\top}w_{y},$
Simultaneously, for OOD queries with feature embedding $z_{\text{OOD}}\in\mathbb{R}^{d}$, we penalize cosine similarity with all class weight vectors, thereby pushing them into a subspace orthogonal to the span of $\{w_{1},w_{2},\dots,w_{C}\}$, where $C$ is the number of clusters as resulted from the previous step. This is achieved through the feature separation loss:
$\mathcal{L}_{\text{Sep}}=\frac{1}{C}\sum_{i=1}^{C}\left|z_{\text{OOD}}^{\top}w_{i}\right|.$
In addition to these, we also include the Outlier Exposure (OE) loss (Hendrycks, Mazeika, and Dietterich 2018) for OOD samples:
$\mathcal{L}_{\text{OE}}(x)=-\frac{1}{C}\sum_{j=1}^{C}\log f_{j}(x),$
where $f_{j}(x)$ denotes the model’s output probability for class $j$.
The total loss function used during fine-tuning is:
$\min_{\theta}\quad&\mathbb{E}_{(x,y)\sim D_{\text{ID}}}\left[\mathcal{L}_{\text{CE}}(x,y)+\alpha\mathcal{L}_{\text{Clu}}\right]\\ &+\mathbb{E}_{x\sim D_{\text{OOD}}^{\text{aux}}}\left[\lambda\mathcal{L}_{\text{OE}}(x)+\beta\mathcal{L}_{\text{Sep}}\right]\end{aligned}$
where $\alpha$, $\beta$, and $\lambda$ are hyperparameters, and $\mathcal{L}_{\text{CE}}$ is the usual Cross-Entropy loss.
Testing.
At inference, we compute the model’s output $f(x)$ and extract its final-layer embedding $z$. Following Wu et al. (2024), we define the score:
$S(x)=\;&\underbrace{\max_{i}\frac{e^{f_{i}(x)}}{\sum_{j=1}^{C}e^{f_{j}(x)}}}_{\text{MSP score}}+\underbrace{\frac{1}{C}\sum_{i=1}^{C}\left|z^{\top}w_{i}\right|}_{\text{Feature separation score}}\end{aligned}$
A higher score indicates a greater likelihood of being ID. To set a threshold $\tau$, we compute the score distribution for train data and set $\tau$ such that the true positive rate of train samples is at 95%.
A new question is classified as:
$\text{ID}\quad\text{if }S(x)\geq\tau,\quad\text{OOD otherwise}.$
This threshold-based rule allows reliable discrimination between ID and OOD queries based on both the model’s output and its latent feature space.
RAG Evaluation
@@
In the second part of our study, we aim to evaluate a simplistic RAG system’s responses in terms of two dimensions: relevance and correctness. Our RAG system follows the approach of Lewis et al. (2020). Initially, a BERT-based bi-encoder model is utilized to compute embeddings for the ID queries $q_{1},\ldots,q_{n}$ offline. These embeddings are then stored within the Retriever component for efficient access during inference. [^3] As our approach is standard, further details are in Appendix.
@@
We conduct human evaluation to assess relevance and correctness. In parallel, we utilized a Large Language Model (LLM-as-a-judge Zheng et al. (2023)) to independently assess the relevance and correctness of each pair, allowing us to compare human vs. LLM-generated evaluations. The templates for relevance and correctness judgments are in Appendix.
Experiment
Data
@@
Our main COVID-19 dataset is from the chatbot logs of a deployed dialog system (VIRA) for COVID-19 vaccine information (Gretz et al. 2023) as well as the KB. We additionally include a 4chan attack that is not available in the default VIRA logs. We note that for the results of Table 2, we extracted a subset of 201 samples from 4chan set, where we manually labeled them as ID or OOD. The dataset for the Substance Use (SU) domain consists of 629 question-answer pairs. This KB includes several domains, including various legal and illegal substances, mental health, treatment, and recovery (see Appendix for exact sources). Furthermore, we use the standard dataset from MS MARCO and StackExchange (Bajaj et al. 2016; Team 2021).
In addition to established datasets, we use a novel synthetic LLM-generated dataset. The existing literature documents generation of datasets through LLMs for applications such as toxicity detection (Hartvigsen et al. 2022; Kruschwitz and Schmidhuber 2024). For this study, we employed GPT-4o to generate the dataset. The inputs of the LLM that affected the generated output are the utilized prompt ($P$), the COVID-19 dataset queries ($Q$), and the chatbot’s response to them ($R$), as illustrated in the formula: $o=f(P,Q,R)$ where $f$ is the generation model (See Appendix for the complete prompt). Table 5 presents selected examples of the generated queries.
@@
In our second study, we extracted 150 ID and 150 OOD samples. The ID queries were generated as follows: we randomly selected 150 samples from our COVID-19 dataset and then applied a rephrasing task using GPT-4o (detailed prompt in Appendix) to transform them into natural user-like questions. For the OOD queries, we randomly selected 150 samples from the larger 4chan dataset, ensuring no duplicates were included. To analyze the distribution of these queries, we visualize histograms of their distances from our KB in Fig. 2. The distance calculation was based on a semantic space where COVID-19 samples were treated as positive instances, while the rest 4chan samples (excluding those in the test set) were considered negative. The PCs were ranked using the p-value criterion, and the optimal number of PCs (p = 15) was selected based on our experimental results to compute the semantic distance (defined as the minimum distance from any sample in the KB).
A smaller-scale study was also conducted for the SU domain. The process for generating the final samples for annotation followed the same steps as in COVID-19 case, except that 75 ID and 75 OOD samples were chosen.

### v2

# Knowing When Not to Answer: Lightweight KB-Aligned OOD Detection for Safe RAG

@@ Abstract
Retrieval-Augmented Generation (RAG) systems are increasingly deployed in high-stakes domains, where safety depends not only on how a system answers, but also on whether a query should be answered given a knowledge base (KB). Out-of-domain (OOD) queries can cause dense retrieval to surface weakly related context and lead the generator to produce fluent but unjustified responses. We study lightweight, KB-aligned OOD detection as an always-on gate for RAG systems. Our approach applies PCA to KB embeddings and scores queries in a compact subspace selected either by explained-variance retention (EVR) or by a separability-driven $t$-test ranking. We evaluate geometric semantic-search rules and lightweight classifiers across 16 domains, including high-stakes COVID-19 and Substance Use KBs, and stress-test robustness using both LLM-generated attacks and an in-the-wild 4chan attack. We find that low-dimensional detectors achieve competitive OOD performance while being faster, cheaper, and more interpretable than prompted LLM-based judges. Finally, human and LLM-based evaluations show that OOD queries primarily degrade the relevance of RAG outputs, showing the need for efficient external OOD detection to maintain safe, in-scope behavior. [^1] https://github.com/toastedqu/rag_safety_pca
Knowing When Not to Answer: Lightweight KB-Aligned
OOD Detection for Safe RAG
Ilias Triantafyllopoulos1, Renyi Qu4, Salvatore Giorgi2, Brenda Curtis2, Lyle H. Ungar3, João Sedoc1 1New York University, 2National Institute on Drug Abuse, 3University of Pennsylvania, 4Vectara, Inc.
1 Introduction
@@
In high-stakes domains, the accuracy and domain relevance of responses provided by Retrieval-Augmented Generation (RAG) systems are critical for ensuring safety and reliability. One significant challenge these systems face is the detection and handling of out-of-domain (OOD) queries, which can impair performance and safety. For instance, in the medical field, a RAG system for clinical decision support must accurately discern relevant medical information (Giorgi et al., 2024). A failure to do so (such as treating an OOD query about a rare medical condition as if it were in-domain (ID)) could result in incorrect medical advice, posing adverse health outcomes. An example is illustrated in Fig. 1 for COVID-19 that shows how an OOD query (right side) can bypass a system’s safeguards and produce a potentially malicious response.
For Knowledge Base (KB)-backed assistants, “safety” decomposes into two orthogonal dimensions (Fig. 2): (A) Answerability / KB alignment—whether the fixed KB can support a grounded answer to the query (our OOD notion), and (B) Intent / policy risk—whether the request is permissible to answer even if it is KB-supported. This paper targets Axis A via lightweight, KB-aligned OOD detection; Axis B requires separate policy filters / content moderation. Accordingly, adversarial or toxic queries in our evaluation are treated as stress tests for Axis A (KB-unsupported queries), rather than a complete solution to harmful but in-domain requests.
Figure 2: Two-axis view of “when not to answer” for KB-backed assistants. We focus on the vertical axis.
A key deployment reality is that a KB-backed assistant must decide not only how to answer, but also whether it should answer at all. While classic and modern RAG architectures effectively ground generation for ID queries (Lewis et al., 2020; Guu et al., 2020; Izacard et al., 2023; Ram et al., 2023; Li et al., 2022; Jiang et al., 2023; Cheng et al., 2024; Zakka et al., 2024), they typically assume that user inputs are relevant to the KB and therefore answerable. When this assumption breaks, two safety-critical failures arise: (i) retrieval may surface weakly related context (dense retrieval is brittle under lexical variation and domain shift) and (ii) the generator may still produce fluent but unjustified responses (“hallucinations”) (Barnett et al., 2024; Reichman and Heck, 2024; Chandrasekaran and Mago, 2021; Bang et al., 2023; Xu et al., 2024; Zhang et al., 2023). As a result, OOD queries can turn an otherwise grounded pipeline into a system that is confident, costly, and wrong.
Practitioners typically address this challenge by relying on built-in LLM guardrails or by adding extra LLM calls to judge whether a query is ID (Peng et al., 2025). However, guardrails primarily target toxic or unsafe content rather than domain irrelevance, offering limited protection against benign but unanswerable queries (Dong et al., 2024). LLM-based “domain judges” also incur substantial latency and API cost, making them impractical as always-on gates. This motivates a complementary goal: a lightweight external OOD detector that is fast, cheap, interpretable, and competitive with LLM-based approaches.
We pursue this goal through a simple principle: domain membership should be testable in a compact representation aligned with the KB. We apply Principal Component Analysis (PCA) to document embeddings and project queries into the resulting subspace, selecting components via either (i) explained variance (EVR) or (ii) a separability-driven statistical test between ID and OOD projections. On this low-dimensional space, we evaluate geometric semantic-search rules and lightweight classifiers. Beyond detection accuracy, we study downstream RAG behavior by comparing an LLM-only pipeline to a two-stage system that abstains based on our detector. Across high-stakes domains (COVID-19 and Substance Use), multi-domain benchmarks, and both LLM-generated and in-the-wild attacks, we show that simple, KB-aligned detectors provide effective protection while substantially reducing inference latency and cost.
Our contributions are as follows:
  • We introduce a KB-aligned, PCA-based OOD detector for RAG and compare variance-based (EVR) and separability-based principal components selection for learning compact, discriminative subspaces
  • We systematically evaluate geometric semantic-search rules and lightweight classifiers across four datasets spanning 16 domains, including real-world high-stakes KBs and both synthetic (LLM-generated) and real attack data
  • We link OOD detection to end-to-end RAG behavior via human and LLM-as-a-Judge evaluation, showing that external OOD detection is necessary to preserve response relevance when guardrails alone are insufficient, while offering substantial practical benefits such as interpretability
2 Related Work
2.1 Safety Concerns in RAG systems
@@
RAG systems rely on the retrieval of relevant documents to ensure accurate and trustworthy responses Lewis et al. (2020), yet incorrect retrievals can severely degrade the quality of generation Creswell et al. (2022); Barnett et al. (2024). Efforts to improve retrieval include incorporating topical context Ahn et al. (2022), conversation history Shuster et al. (2021), and predictive sentence generation (FLARE) Jiang et al. (2023). Interestingly, unrelated documents sometimes enhance generation, while highly ranked but irrelevant ones can harm it Cuconasu et al. (2024). To mitigate such issues, methods have proposed response skeletons Cai et al. (2019), prompt-based validation Yu et al. (2023), Natural Language Inference (NLI) filtering Yoran et al. (2023), and dynamic reliance on parametric vs. retrieved knowledge Li et al. (2023); Longpre et al. (2021); Mallen et al. (2023). Our work adds to this literature by introducing new methods for selecting when to answer and verifying the effect of out-of-domain questions.
2.2 Adversarial Attacks
@@
Adversarial attacks mislead models through crafted inputs Zhang et al. (2020), with recent work targeting LLMs to produce harmful content Zou et al. (2023). For RAG systems, attacks often involve malicious documents that degrade retrieval or generation Cho et al. (2024); Xue et al. (2024); Shafran et al. (2024). While early attacks required specific trigger queries Zou et al. (2024), newer methods exploit query-agnostic poisoning Chaudhari et al. (2024). Our work mitigates query-based adversarial attacks by detecting when a modified question lacks an answer in the database. This work focuses on "attacks" that are an outcome of real-world questions.
3 Methods
3.1 Out-Of-Domain Detection
Given a user query $q$ and a document dataset $\mathcal{D}=\{d_{1},\cdots,d_{n}\}$, where $n$ is the total number of documents, the detection of OOD queries aims to predict whether the query is relevant to the document space and therefore answerable by the response generation module, which is typically performed by an LLM. Our method is straightforward. First, we compute the query embedding $\mathbf{e}_{q}$ and the document embeddings $E_{d}=[\mathbf{e}_{d_{1}};\cdots;\mathbf{e}_{d_{n}}]$ using a pretrained BERT-based bi-encoder model. Second, we run PCA on document embeddings to retrieve the top-$k$ principal components (PCs), denoted $PC_{k}=[\mathbf{pc}_{1},\mathbf{pc}_{2},\cdots,\mathbf{pc}_{k}]$, which represent the dominant patterns within the document space and capture the largest variances. After determining the top-$k$ PCs, we further refine the selection to a final set of $m$ PCs using two different criteria:
@@ 3.1 Out-Of-Domain Detection
  • Explained Variance (EVR): In this approach, the final set of PCs consists of the $m$ components with the highest explained variance, where $m\leq k$ (note that this is effectively top-$m$). This ensures that we retain only the components that contribute the most to the variance of the data set.
  • p-values: Here, we project the query embeddings of both ID queries and OOD queries onto the document embeddings. A t-test is conducted for each dimension of the top-$k$ PCs, comparing the ID and OOD query projections. The PCs are then sorted by their $p$-values in ascending order, and the $m$ PCs with the lowest $p$-values are selected. This approach ensures that the retained dimensions are the most effective in distinguishing ID and OOD queries.
@@
Both criteria for selecting $m$ PCs aim to retain the most informative aspects of the embeddings while reducing dimensionality. This step not only preserves the discriminative power of the embeddings but also enhances computational efficiency for subsequent tasks.
Third, we project the query embedding $\mathbf{e}_{q}$ onto this reduced space to obtain a transformed query embedding $\mathbf{e}_{q}^{\prime}$. This transformation is crucial as it allows the query’s position relative to the PCs of the document space to be quantified, enabling a more accurate assessment of its relevance. Specifically, we use the projection formula $\mathbf{e}_{q}^{\prime}=\mathbf{e}_{q}PC_{m}^{T}$, where $PC_{m}^{T}$ is the transpose of the matrix containing the top-$m$ PCs.
@@
The semantic-search algorithms operate by mapping the query embeddings from the training set into a $m$-dimensional space derived from the PC selection process. During inference, these algorithms employ distinct geometric criteria to make a classification decision for a test query $u$ with projected embedding $\mathbf{e}_{u}^{\prime}$. The three algorithms are as follows:
@@
In addition to the semantic-search algorithms, we leverage three simple yet effective machine learning models. These models are trained on the entire training set, which includes both ID and OOD queries, for a binary classification task. During inference, the algorithms classify the test query $u$ with its projected embedding $\mathbf{e}_{u}$ into one of the two classes. We use the following models: a Logistic Regression (LogReg) (Kleinbaum et al., 2002), Support Vector Machines (SVM) (Hearst et al., 1998), and Gaussian Mixture Models (GMM) (Reynolds and others, 2009).
3.2 RAG Evaluation
@@
In the second part of our study, we aim to evaluate a simplistic RAG system’s responses in terms of two dimensions: relevance and correctness. Our RAG system follows the approach of Lewis et al. (2020). Initially, a BERT-based bi-encoder model is utilized to compute embeddings for the ID queries $q_{1},\ldots,q_{n}$ offline. These embeddings are then stored within the Retriever component for efficient access during inference. [^2] As our approach is standard, further details are in Appendix E.
@@
We conduct human evaluation to assess relevance and correctness. In parallel, we utilized a Large Language Model (LLM-as-a-judge Zheng et al. (2023)) to independently assess the relevance and correctness of each pair, allowing us to compare human vs. LLM-generated evaluations. The templates for relevance and correctness judgments are in Appendix A.
This simplified setup serves two purposes. First, it reflects a common deployment pattern in which dense retrieval is combined with a general-purpose LLM, making the evaluation representative of real-world systems. Second, it allows us to study the effect of OOD detection independently of architectural complexity. This evaluation is particularly important in high-stakes settings, where a system that produces fluent but irrelevant responses can be more harmful than one that abstains.
4 Experiments
4.1 Data
@@
Our main COVID-19 dataset is from the chatbot logs of a deployed dialog system (VIRA) for COVID-19 vaccine information Gretz et al. (2023). We additionally include a 4chan attack that is not available in the default VIRA logs, but comes from real-world attacks that occurred against the chatbot. We note that for Table 1 results, we extracted a subset of 201 samples from 4chan set, where we manually labeled them as ID or OOD. The dataset for the Substance Use (SU) domain consists of 629 question-answer pairs. This KB includes several topics, including various legal and illegal substances, mental health, treatment, and recovery (see Appendix B for exact sources). Furthermore, we use the standard MS MARCO and StackExchange datasets Bajaj et al. (2016); Team (2021).
In addition to established datasets, we construct a synthetic LLM-generated dataset. Prior work has used LLMs to generate datasets for tasks such as toxicity detection (Hartvigsen et al., 2022; Kruschwitz and Schmidhuber, 2024). We employ GPT-4o to generate queries conditioned on the prompt ($P$), COVID-19 dataset queries ($Q$), and the chatbot’s responses ($R$), formalized as $o=f(P,Q,R)$, where $f$ denotes the generation model (see Appendix A for the full prompt). Table 6 shows representative generated queries.
@@
In our second study, we construct a balanced evaluation set of 150 ID and 150 OOD samples. ID queries are created by randomly selecting 150 COVID-19 queries and rephrasing them into natural, user-like questions using GPT-4o (see Appendix A). OOD queries are sampled from the larger 4chan dataset, with duplicates removed. To analyze their distribution, we visualize histograms of each query’s semantic distance from the KB (Fig. 3). Distances are computed in a semantic space where COVID-19 queries are treated as ID, and the remaining 4chan queries (excluding the test set) as OOD. PCs are ranked using the p-value criterion, and the optimal dimensionality ($p=15$) is selected based on experimental results; semantic distance is defined as the minimum distance to any KB sample. We additionally conduct a smaller-scale study for the SU domain following the same procedure, using 75 ID and 75 OOD samples.